OpenStudy (anonymous):

Caley has a bag that contains 9 blue marbles and 13 yellow marbles. She selects a marble at random, and then, without replacing the first one, selects another marble at random. What is the probability that Caley selects a blue marble and then a yellow marble? Round your answer to the nearest percent. P(blue and yellow) ≈ %

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

how many blue marbles and how many total?

OpenStudy (anonymous):

9 & 22

Mehek (mehek14):

so \(\dfrac{9}{22}\) now when you take the blue out and don't put it back, the total number of outcome changes to 21

Mehek (mehek14):

so how many yellow marbles?

OpenStudy (anonymous):

13?

Mehek (mehek14):

ok so you have \(\dfrac{9}{22} * \dfrac{13}{21}\)

Mehek (mehek14):

now multiple them

OpenStudy (anonymous):